# Exploring Self-Interacting Dark Matter Halos with Diverse Baryonic Distributions: A Parametric Approach

### v1

@@ VII.1 Probing SIDM through sampled halos
For the study here, we analyze the CDM and two variants of SIDM models. The first SIDM model employs a constant cross section per mass, set at $\sigma/m=10~{}\rm cm^{2}/g$. The second SIDM model uses a velocity-dependent cross section, modeled assuming Rutherford-like scatterings, as described by
@@
where $\sigma_{0}/m=100~{}\rm cm^{2}/g$ and $w=100~{}\rm km/s$.
Our findings are depicted in Fig. 8, organized from left to right, showcasing results for the CDM, constant cross section SIDM, and velocity-dependent SIDM models. Each model is examined across three rows of figures to elucidate the effects of baryons: the top row presents the dark matter-only scenario, while the middle and bottom rows explore the combined influence of dark matter and baryons. Specifically, the middle row visualizes $V_{\rm max}$ and $R_{\rm max}$ for the dark matter component alone, and the bottom row incorporates the baryonic contributions.
@@
For the SIDM models, results are color-coded based on the halos’ gravothermal phases at $z=0$. In both models, baryonic effects are pronounced: data distribution spreads wider, and halos hosting massive galaxies can exhibit core-collapsing phases characterized by decreased $R_{\rm max,DM}$ and slightly increased $V_{\rm max,DM}$. We also explore novel correlations between SIDM model parameters and halo mass scales. In the constant cross section model ($\sigma/m=10~{}\rm cm^{2}/g$), a significant number of halos are in core-collapsing phases at $V_{\rm max}\gtrsim 300~{}\rm km/s$. Conversely, the velocity-dependent model shows a higher incidence of core collapse at lower velocities ($V_{\rm max}\lesssim 60~{}\rm km/s$).
@@ VIII Conclusion and discussion
Crucially, the assumption that each SIDM halo is in a gravothermal state defined by $\tau$ operates independently of the parametric model. This assumption allows for the use of controlled N-body simulations to accurately model any halo’s state within a cosmological simulation, as long as their $\tau$ phases match. For dark matter-only halos, the validity of this assumption is corroborated by comparisons between the model and simulations, as detailed in Refs. Yang et al. (2024a) and Yang et al. (2024b), the latter of which examines the relative differences on a case-by-case basis. In scenarios involving baryons, however, verification via N-body simulations remains to be shown. Nonetheless, as shown in Appendix B, the reliability of this assumption supports theoretical explorations. Conducting dedicated N-body tests of this assumption would facilitate obtaining accurate theoretical predictions in future studies.
@@ Acknowledgments
I would like to thank Hai-Bo Yu, Ethan O. Nadler, Yi-Ming Zhong, Shin’ichiro Ando, and Demao Kong, for their helpful discussions. Special appreciation goes to Hai-Bo Yu for helpful comments on fitting rotation curves and SIDM’s feedback on baryons. I also acknowledge the IllustrisTNG collaborations for making their simulation data publicly available. This work was supported by the John Templeton Foundation under Grant ID #61884, and the U.S. Department of Energy under grant No. de-sc0008541. The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton Foundation.

### v2

@@ VII.1 Probing SIDM through sampled halos
In Fig. 8, we present the parametric model predictions for three scenarios: CDM (left column), a constant SIDM cross section of $\sigma/m=10,\rm cm^{2}/g$ (middle column), and a velocity-dependent SIDM cross section (right column) given by
@@
where $\sigma_{0}/m=100~{}\rm cm^{2}/g$ and $w=100~{}\rm km/s$. Each column displays results across three cases: dark matter-only (top row), dark matter with baryon feedback (middle row), and full matter including baryons (bottom row).
@@
In the SIDM models, color-coding reflects halos’ gravothermal phases at $z=0$. Here, baryonic effects are pronounced, widening the data spread and, in massive galaxies, inducing core-collapsing phases with decreased $R_{\rm max,DM}$ and slightly increased $V_{\rm max,DM}$. The constant cross section model ($\sigma/m=10~{}\rm cm^{2}/g$) shows core-collapsing halos predominantly appear at high $V_{\rm max}\gtrsim 300~{}\rm km/s$, whereas the velocity-dependent model displays more core collapse at lower velocities ($V_{\rm max}\lesssim 60~{}\rm km/s$).
The middle row of Fig. 8 presents results for the dark matter component alone, illustrating the effect of baryons on dark matter content, with and without SIDM, to address theoretical interests. The bottom row presents more realistic outcomes, visualizing $V_{\rm max}$ and $R_{\rm max}$ with baryonic contributions included. Comparing these with the middle panels reveals how baryons diminish SIDM signatures, especially around $V_{\rm max}\approx 200~{}\rm km/s$, where the baryon-to-dark-matter mass ratio peaks. Nevertheless, distinct differences remain across the three cases after accounting for baryons. Reducing observational uncertainty will be crucial to test analogous SIDM scenarios using variables related to rotation curves.
@@ VIII Conclusion and discussion
Crucially, the assumption that each SIDM halo is in a gravothermal state defined by $\tau$ operates independently of the parametric model. This assumption allows for the use of controlled N-body simulations to accurately model any halo’s state within a cosmological simulation, as long as their $\tau$ phases match. For dark matter-only halos, the validity of this assumption is corroborated by comparisons between the model and simulations, as detailed in Refs. Yang et al. (2024a) and Yang:2024uqb, the latter of which examines the relative differences on a case-by-case basis. In scenarios involving baryons, however, verification via N-body simulations remains to be shown. Nonetheless, as shown in Appendix B, the reliability of this assumption supports theoretical explorations. Conducting dedicated N-body tests of this assumption would facilitate obtaining accurate theoretical predictions in future studies.
@@ Acknowledgments
I would like to thank Hai-Bo Yu, Ethan O. Nadler, Yi-Ming Zhong, Shin’ichiro Ando, and Demao Kong, for their helpful discussions. Special appreciation goes to Hai-Bo Yu for helpful comments on fitting rotation curves and SIDM’s feedback on baryons. I also acknowledge the IllustrisTNG collaborations for making their simulation data publicly available. This work was supported in part by the U.S. Department of Energy under grant No. de-sc0008541, the National Key Research and Development Program of China (No. 2022YFF0503304), the Project for Young Scientists in Basic Research of the Chinese Academy of Sciences (No. YSBR-092), and the John Templeton Foundation under grant ID #61884. The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton Foundation.